# The HST-Hyperion Survey: Environmental Imprints on the Stellar-Mass Function at $z\sim 2.5$

### v1

@@ Abstract
Context. Not all galaxies at Cosmic Noon ($2\lesssim z\lesssim 3$) evolve in the same way. Particularly, it remains unclear how and to what extent the local environment – especially the extreme overdensities of protoclusters – affects the stellar mass assembly of its constituent galaxies at high redshifts. The imprint of those early processes is encoded in the galaxy stellar-mass function (SMF); comparing SMFs across environments therefore reveals differences in evolutionary history.
Aims. We present the SMF of the Hyperion proto-supercluster at $z\sim 2.5$ – one of the largest and most massive protostructures in the early universe. This dataset yields the most statistically robust SMF of a single protostructure at $z\gtrsim 2$. By comparing the SMF of the overdense peaks within Hyperion to the coeval field, we begin to answer: How early, and how strongly, does a dense environment tilt the balance in favor of massive galaxies?
Methods. Given Hyperion resides in the field of the Cosmic Evolution Survey (COSMOS), we combine the extensive COSMOS2020 photometric catalog with ground-based spectroscopy and new grism spectroscopy from the Hubble Space Telescope (HST). The structure of Hyperion is defined based on a three-dimensional overdensity map, allowing us to place galaxies into either (i) the highly-overdense peaks of Hyperion, (ii) the less-overdense outskirts of Hyperion, or (iii) a coeval field. We perform 100 Monte Carlo realizations of the data to propagate redshift and stellar mass uncertainties, refitting galaxy properties in each realization. After constructing SMFs for the outskirts and peaks of Hyperion, we normalize them to that of the field to highlight differences in the underlying shape of the SMFs.
@@
Conclusions. Environmental effects that govern stellar mass growth are already well-established by $z\sim 2.5$: the densest regions of Hyperion are host to galaxies which have already experienced accelerated stellar mass growth. Furthermore, this impact is largely masked in the total SMF of Hyperion, highlighting the necessity of deep spectroscopic surveys when uncovering environmental trends at high redshifts. These findings imply that high redshift protostructures begin sculpting the high-mass end of the SMF well before the epoch when local clusters experience widespread quenching, and may provide the appropriate laboratories for producing the elevated star formation observed at Cosmic Noon.
@@ 1 Introduction
It has long been understood that different galaxies can experience dramatically different evolutionary pathways, resulting in distinct physical properties. This is particularly evident at low and intermediate redshifts ($z\lesssim 1.5$) where galaxies residing in dense cores of clusters at low redshifts tend to display higher stellar masses (e.g., Kauffmann et al. 2004; Baldry et al. 2006; van der Burg et al. 2020), lower star formation rates (SFRs; e.g., Gómez et al. 2003; von der Linden et al. 2010; Muzzin et al. 2012; Tomczak et al. 2019; Old et al. 2020), and redder colors (e.g., Peng et al. 2010; Lemaux et al. 2012; Nantais et al. 2016; Lemaux et al. 2019; van der Burg et al. 2020) compared to their counterparts in the coeval field. Under the assumption that these two apparently different populations of galaxies formed with similar initial properties, this bimodality implies the existence of environmentally-dependent processes which have profound effects on the way galaxies evolve. However, the exact nature of the dominant processes responsible for this differential evolution and the time period in which they act on remains elusive.
By Cosmic Noon ($2\lesssim z\lesssim 3$), the picture seemingly flips: several studies now suggest that overdense environments are conducive to higher star formation rate densities (SFRDs) than the field (e.g., Lemaux et al. 2022; Staab et al. 2024). This is a stark contrast to the local universe, and suggests a more complex redshift-dependence of the environmentally-dependent processes driving galaxy evolution. Taken together with local trends, it seems as though the same environments which eventually suppress stellar mass growth in the local universe may in fact accelerate it in the early universe.
Given that the key mechanisms spurring galaxy evolution likely take place over prolonged timescales, catching them in the act for any given galaxy is difficult. However, their integrated impact is encoded in the stellar mass function (SMF) – the number density of galaxies as a function of stellar mass. In the local universe, various studies focused on the SMF have contributed to a better understanding of the processes driving the observed bimodality of galaxies. In particular, Peng et al. (2010) developed an empirical model in which environmental quenching truncates star formation irrespective of mass, while mass quenching limits growth in the most massive systems regardless of environment. At intermediate redshifts, studies such as Tomczak et al. (2017) find the SMF depends strongly on environment, with overdense environments hosting an excess of higher stellar mass compared to lower stellar mass galaxies, relative to the field.
@@ 2.1 COSMOS2020 Photometry
The photometric measurements used in this work are taken from the COSMOS2020 catalogs (Weaver et al. 2022). These catalogs contain extensive data for $\sim\!1.7$ million objects spanning $\sim\!2\,\,\rm deg^{2}$, including imaging in up to 40 bandpasses and physical properties estimated from two Spectral Energy Distribution (SED) fitting codes. For this work, we specifically use COSMOS2020 Classic Catalog v2.0 and physical parameters derived with LePhare [^1] LePhare’s website: http://www.cfht.hawaii.edu/ arnouts/lephare.html (Arnouts et al. 1999; Ilbert et al. 2006).
@@
In the work that follows, we use flux densities for all bands where available, except for the HST/ACS F814W band and Spitzer/IRAC channels 3 and 4. Specifically, for a given band XXXX, we use the flux-density reported under XXXX_FLUX_APER3 in the COSMOS2020 catalog and its associated error, which was derived using a 33"-aperture in SExtractor (Bertin & Arnouts 1996). For each source, an aperture-to-total flux correction using total_off3 is applied to the flux-density in all bands taken from ground-based observations. Finally, a foreground attenuation correction is applied using the extinction term EBV_MW, along with a multiplicative band-dependent factor, as described in Laigle et al. (2016).
We impose additional selection criteria to define our final sample of COSMOS2020 photometric sources. We limit our analysis to objects labeled as galaxies or X-ray sources in the COSMOS2020 catalog (lp_type = 0 or 2), eliminating 39,653 sources. We further restrict the data to objects within the R.A. and decl. range of Hyperion, as determined in Cucciati et al. (2018) (see also Section 3). After requiring $\rm 149.60^{\circ}\leq R.A.\leq 150.52^{\circ}$ and $\rm 1.74^{\circ}\leq decl.\leq 2.73^{\circ}$, we retain 498,273 objects. Finally, to ensure completeness, we impose a magnitude cut in the IRAC bands, requiring either $3.6]\leq 24.8$ or $4.5]\leq 24.8$, leaving 220,356 sources. We make no cut on redshift at this point as a Monte Carlo routine will be used later on with the reported COSMOS2020 redshift PDFs (see Section 4). The redshift distribution of the remaining photometric sources is plotted in Figure 1.
Figure 1: Redshift Distribution of Photometry – The redshift distribution of the 220,356 objects used in this study from the COSMOS2020 catalog after applying the cuts on object type, R.A., decl., and IRAC magnitude described in Section 2.1. The photometric redshifts are taken from lp_zPDF in the COSMOS2020 catalog. Note that these are the objects that may have their redshift PDFs sampled during the Monte Carlo process (see Section 4.1). We additionally mark the approximate boundaries of Hyperion in redshift-space.
@@ 2.2.1 Wide-Field Surveys
There have been a number of spectroscopic surveys which have added to the wealth of data in the COSMOS field. The zCOSMOS survey (Lilly et al. 2007) was conducted using the Visible Multi Object Spectrograph (VIMOS) on the Very Large Telescope (VLT), culminating in 600 hours of observations. It consists of two components: zCOSMOS-bright, a magnitude-limited sample spanning $0.1<z<1.2$, and zCOSMOS-deep, a sample of $\sim 10,000$ color-selected galaxies in the redshift range $1.4<z<3.0$ within the central $\rm 1\,\deg^{2}$ of the COSMOS field. Also taken on VIMOS was the VIMOS Ultra Deep Survey (VUDS; Le Fèvre et al. 2015), which includes spectroscopy of $\sim 10,000$ galaxies ($\sim 5000$ in the COSMOS field) with $i<25$, spanning a redshift range of $2<z<6$. Additionally, the DEIMOS 10k survey (Hasinger et al. 2018) using the Deep Imaging Multi-Object Spectrograph (DEIMOS) (Faber et al. 2003) on the Keck II telescope took spectra of over $10,000$ objects with $i<23$ in the redshift range of $0<z<6$, though the redshift distribution was heavily weighted to lower redshifts (see top panel of Figure 2).
To ensure the reliability of redshift measurements across these surveys, a standardized system of quality flags is used. zCOSMOS and VUDS follow a similar system for assigning quality flags to the spectroscopic redshifts, based on Section 6.5 of Le Fèvre et al. (2005). The redshift quality flags from the DEIMOS 10k survey were subsequently conformed to fit the same system (Lemaux et al. 2022). Briefly, objects are assigned a flag from 0, 1, 2, 3, 4, or 9 to indicate the confidence of the redshift measurement. A complete discussion of the flag assignment and redshift reliability assessment is presented in Appendix A of Lemaux et al. (2022), but the redshift qualities can be summarized as follows:
Flags 3/4: High‐confidence ($\sim 99.3\%$ reliable).
Flags 2/9: Moderate‐confidence ($\sim 70\%$ reliable).
@@
Each quality flag can be prepended with additional flags in order to indicate if the object is peculiar in some way. For example, a quality flag of 13 would indicate a broadline AGN (per the prepended 1) with a secure redshift. As we are not interested in the specifics of these objects in this respect for this analysis, we ignore the prepended flags and only consider spectra which were assigned quality flags ending with 2, 3, 4, or 9 regardless of what is prepended.
In addition to the quality flag criteria, we also require the sources have COSMOS2020 photometry so that physical properties can be derived via SED fitting (Section 4.2), and are identified as either galaxies or X-ray sources in COSMOS2020. Furthermore, we only consider objects falling around Hyperion with $\rm 149.60^{\circ}\leq R.A.\leq 150.52^{\circ}$ and $\rm 1.74^{\circ}\leq decl.\leq 2.73^{\circ}$, and which have either $3.6]\leq 24.8$ or $4.5]\leq 24.8$ (from COSMOS2020). After applying these cuts, we are left with 11,368 objects from the zCOSMOS survey, 4,410 objects from the DEIMOS 10k survey, and 2,129 objects from the VUDS survey. The redshift distribution of these objects is shown in the top panel of Figure 2.
@@ 2.2.2 Targeted Surveys
We also draw on spectra from the C3VO program – the Charting Cluster Construction with VUDS (Le Fèvre et al. 2015) and ORELSE (Lubin et al. 2009) survey – which uses the DEIMOS (Faber et al. 2003) and Multi-Object Spectrometer For Infra-Red Exploration (MOSFIRE; McLean et al. 2012) instruments on Keck to follow up protocluster candidates selected with the Voronoi Monte Carlo (VMC) method (Lemaux et al. 2022). Across the fields of the Canada-France-Hawai’i Telescope Legacy Survey D1 (CFHTLS-D1), the Extended Chandra Deep Field South (ECDFS), and COSMOS, C3VO has obtained $\sim\!2000$ high-quality redshifts, yielding detailed maps of structures such as PCl J1000+0200 ($z\sim 2.90$; Cucciati et al. 2014), PCl J0227-0421 ($z\sim 3.31$; Lemaux et al. 2014; Shen et al. 2021), Elentári ($z\sim 3.33$; McConachie et al. 2022; Forrest et al. 2023), Smruti ($z\sim 3.47$; Forrest et al. 2017; Shah et al. 2024a), Taralay ($z\sim 4.57$; Lemaux et al. 2018; Staab et al. 2024), and – central to this work – Hyperion.
@@
Additionally, we include data from the Massive Ancient Galaxies at $z>3$ Near-infrared Survey (MAGAZ3NE; Forrest et al. 2024a), a spectroscopic follow-up program targeting ultra-massive galaxies ($\log_{10}(M_{*}/M_{\odot})\gtrsim 11$ at $z>3$) using Keck/MOSFIRE in multiple fields, including COSMOS. While the galaxies from this survey do not lie within Hyperion, their spectroscopic redshifts provide valuable constraints on photometric redshift outliers. A total of 37 galaxies from MAGAZ3NE meet our selection criteria.
Finally, we incorporate spectra from previous studies that targeted individual peaks of Hyperion. Specifically, we include four spectra from Casey et al. (2015), three from Chiang et al. (2015), and one from Diener et al. (2015), all of which meet the criteria outlined above. Further details regarding observations and data reduction can be found in the respective publications. The redshift distribution of galaxies from the various targeted spectroscopic surveys is shown in the middle panel of Figure 2.
Figure 2: Spectroscopic and HST/Grism Targets – The distribution of spectroscopic redshifts used in this study from (Top) wide-field spectroscopic surveys (section 2.2.1), (Middle) targeted spectroscopic surveys (section 2.2.2), and (Bottom) the HST-Hyperion survey (section 2.3). For each survey, we show only the galaxies which meet the imposed criteria on R.A., decl., and IRAC magnitude described in their respective sections. Ground-based spectroscopic redshifts have all have reliabilities of $\gtrsim 70\%$, whereas HST-grism redshifts all have reliabilities $\gtrsim 56.7\%$. The number of spectra used from each survey is reported in the legend next to the survey name. In each plot, we additionally mark the approximate boundaries of Hyperion in redshift-space.

### v2

# The HST-Hyperion survey: Environmental imprints on the stellar mass function at $z\sim 2.5$

@@ Abstract
Context. Not all galaxies at cosmic noon ($2\lesssim z\lesssim 3$) evolve in the same way. Particularly, it remains unclear how and to what extent the local environment – especially the extreme overdensities of protoclusters – affects the stellar mass assembly of its constituent galaxies at high redshifts. The imprint of these early processes is encoded in the galaxy stellar-mass function (SMF); comparing SMFs across environments therefore reveals differences in evolutionary history.
Aims. We present the SMF of the Hyperion proto-supercluster at $z\sim 2.5$, one of the largest and most massive protostructures in the early universe. This dataset yields the most statistically robust SMF of a single protostructure at $z\gtrsim 2$. By comparing the SMF of the overdense peaks within Hyperion to the coeval field, we begin to answer the question of how early, and how strongly, a dense environment tilts the balance in favor of massive galaxies
Methods. Given that Hyperion resides in the field of the Cosmic Evolution Survey (COSMOS), we combined the extensive COSMOS2020 photometric catalog with ground-based spectroscopy and new grism spectroscopy from the Hubble Space Telescope (HST). The structure of Hyperion is defined based on a three-dimensional overdensity map, allowing us to place galaxies into (i) the highly overdense peaks of Hyperion, (ii) the less-overdense outskirts of Hyperion, or (iii) a coeval field. We performed 100 Monte Carlo realizations of the data to propagate redshift and stellar mass uncertainties, refitting galaxy properties in each realization. After constructing SMFs for the outskirts and peaks of Hyperion, we normalized them to that of the field to highlight differences in the underlying shape of the SMFs.
@@
Conclusions. Environmental effects that govern stellar mass growth are already well established by $z\sim 2.5$. The densest regions of Hyperion host galaxies that have already experienced accelerated stellar mass growth. Furthermore, this impact is largely masked in the total SMF of Hyperion, highlighting the necessity of deep spectroscopic surveys when uncovering environmental trends at high redshifts. These findings imply that high-redshift protostructures begin sculpting the high-mass end of the SMF well before the epoch when local clusters experience widespread quenching, and may provide the appropriate laboratories for producing the elevated star formation observed at cosmic noon.
@@ 1 Introduction
It has long been understood that different galaxies can experience dramatically different evolutionary pathways, thus resulting in distinct physical properties. This is particularly evident at low and intermediate redshifts ($z\lesssim 1.5$) where galaxies residing in dense cores of clusters at low redshifts tend to display higher stellar masses (e.g., Kauffmann et al. 2004; Baldry et al. 2006; van der Burg et al. 2020), lower star formation rates (SFRs; e.g., Gómez et al. 2003; von der Linden et al. 2010; Muzzin et al. 2012; Tomczak et al. 2019; Old et al. 2020), and redder colors (e.g., Peng et al. 2010; Lemaux et al. 2012; Nantais et al. 2016; Lemaux et al. 2019; van der Burg et al. 2020) compared to their counterparts in the coeval field. Under the assumption that these two apparently different populations of galaxies formed with similar initial properties, this bimodality implies the existence of environmentally dependent processes that have profound effects on the way galaxies evolve. However, the exact nature of the dominant processes responsible for this differential evolution and the time period in which they act remain elusive.
By cosmic noon ($2\lesssim z\lesssim 3$), the picture seemingly flips. Several studies now suggest that overdense environments are conducive to higher star formation rate densities (SFRDs) than the field (e.g., Lemaux et al. 2022; Staab et al. 2024). This is a stark contrast to the local universe, and suggests a more complex redshift-dependence of the environmentally dependent processes driving galaxy evolution. Taken together with local trends, it seems as though the same environments that eventually suppress stellar mass growth in the local universe may in fact accelerate it in the early universe.
Given that the key mechanisms spurring galaxy evolution likely take place over prolonged timescales, catching them in the act for any given galaxy is difficult. However, their integrated impact is encoded in the stellar mass function (SMF), i.e., the number density of galaxies as a function of stellar mass. In the local universe, various studies focused on the SMF have contributed to a better understanding of the processes driving the observed bimodality of galaxies. In particular, Peng et al. (2010) developed an empirical model in which environmental quenching truncates star formation irrespective of mass, while mass quenching limits growth in the most massive systems regardless of environment. At intermediate redshifts, studies such as Tomczak et al. (2017) find the SMF depends strongly on environment: overdense environments host an excess of higher stellar mass compared to lower stellar mass galaxies, relative to the field.
@@ 2.1 COSMOS2020 photometry
The photometric measurements used in this work were taken from the COSMOS2020 catalogs (Weaver et al. 2022). These catalogs contain extensive data for $\sim\!1.7$ million objects spanning $\sim\!2\,\,\rm deg^{2}$, including imaging in up to 40 bandpasses and physical properties estimated from two spectral energy distribution (SED) fitting codes. For this work, we specifically use COSMOS2020 Classic Catalog v2.0 and physical parameters derived with LePhare [^1] LePhare’s website: http://www.cfht.hawaii.edu/ arnouts/lephare.html (Arnouts et al. 1999; Ilbert et al. 2006).
@@
In this work we used flux densities for all bands where available, except for the HST/ACS F814W band and Spitzer/IRAC channels 3 and 4. Specifically, for a given band XXXX, we used the flux-density reported under XXXX_FLUX_APER3 in the COSMOS2020 catalog and its associated error, which was derived using a 33" aperture in SExtractor (Bertin and Arnouts 1996). For each source, an aperture-to-total flux correction using total_off3 was applied to the flux density in all bands taken from ground-based observations. Finally, a foreground attenuation correction was applied using the extinction term EBV_MW, along with a multiplicative band-dependent factor, as described in Laigle et al. (2016).
We imposed additional selection criteria to define our final sample of COSMOS2020 photometric sources. We limited our analysis to objects labeled as galaxies or X-ray sources in the COSMOS2020 catalog (lp_type = 0 or 2), eliminating 39,653 sources. We further restricted the data to objects within the R.A. and Dec.. range of Hyperion, as determined in Cucciati et al. (2018) (see also Section 3). After requiring $\rm 149.60^{\circ}\leq R.A.\leq 150.52^{\circ}$ and $\rm 1.74^{\circ}\leq Dec.\leq 2.73^{\circ}$, we retained 498,273 objects. Finally, to ensure completeness, we imposed a magnitude cut in the IRAC bands, requiring either $3.6]\leq 24.8$ or $4.5]\leq 24.8$, leaving 220,356 sources. We made no cut on redshift at this point, as a Monte Carlo routine was used later on with the reported COSMOS2020 redshift PDFs (see Section 4). The redshift distribution of the remaining photometric sources is plotted in Figure 1.
Figure 1: The redshift distribution of the 220,356 objects used in this study from the COSMOS2020 catalog after applying the cuts on object type, R.A., Dec., and IRAC magnitude described in Section 2.1. The photometric redshifts were taken from lp_zPDF in the COSMOS2020 catalog. We note that these are the objects that may have their redshift PDFs sampled during the Monte Carlo process (see Section 4.1). We additionally mark the approximate boundaries of Hyperion in redshift space.
@@ 2.2.1 Wide-field surveys
There have been a number of spectroscopic surveys which have added to the wealth of data in the COSMOS field. The zCOSMOS survey (Lilly et al. 2007) was conducted using the Visible Multi Object Spectrograph (VIMOS) on the Very Large Telescope (VLT), culminating in 600 hours of observations. It consisted of two components: zCOSMOS-bright, a magnitude-limited sample spanning $0.1<z<1.2$, and zCOSMOS-deep, a sample of $\sim 10,000$ color-selected galaxies in the redshift range $1.4<z<3.0$ within the central $\rm 1\,\deg^{2}$ of the COSMOS field. Also taken on VIMOS was the VIMOS Ultra Deep Survey (VUDS; Le Fèvre et al. 2015), which included spectroscopy of $\sim 10,000$ galaxies ($\sim 5000$ in the COSMOS field) with $i<25$, spanning a redshift range of $2<z<6$. Additionally, the DEIMOS 10k survey (Hasinger et al. 2018) using the Deep Imaging Multi-Object Spectrograph (DEIMOS) (Faber et al. 2003) on the Keck II telescope took spectra of over $10,000$ objects with $i<23$ in the redshift range of $0<z<6$, though the redshift distribution was heavily weighted to lower redshifts (see top panel of Figure 2).
To ensure the reliability of redshift measurements across these surveys, a standardized system of quality flags was used. zCOSMOS and VUDS followed a similar system for assigning quality flags to the spectroscopic redshifts, based on Section 6.5 of Le Fèvre et al. (2005). The redshift quality flags from the DEIMOS 10k survey were subsequently conformed to fit the same system (Lemaux et al. 2022). Briefly, objects were assigned a flag from 0, 1, 2, 3, 4, or 9 to indicate the confidence of the redshift measurement. A complete discussion of the flag assignment and redshift reliability assessment is presented in Appendix A of Lemaux et al. (2022), but the redshift qualities can be summarized as follows:
Flags 3/4: High confidence ($\sim 99.3\%$ reliable).
Flags 2/9: Moderate confidence ($\sim 70\%$ reliable).
@@
Each quality flag can be prepended with additional flags in order to indicate if the object is peculiar in some way. For example, a quality flag of 13 indicates a broadline AGN (per the prepended 1) with a secure redshift. As we are not interested in the specifics of these objects in this respect for this analysis, we ignored the prepended flags and only considered spectra which were assigned quality flags ending with 2, 3, 4, or 9 regardless of what is prepended.
In addition to the quality flag criteria, we also required the sources have COSMOS2020 photometry so that physical properties could be derived via SED fitting (Section 4.2), and are identified as either galaxies or X-ray sources in COSMOS2020. Furthermore, we only considered objects falling around Hyperion with $\rm 149.60^{\circ}\leq R.A.\leq 150.52^{\circ}$ and $\rm 1.74^{\circ}\leq Dec.\leq 2.73^{\circ}$, and which have either $3.6]\leq 24.8$ or $4.5]\leq 24.8$ (from COSMOS2020). After applying these cuts, we were left with 11,368 objects from the zCOSMOS survey, 4,410 objects from the DEIMOS 10k survey, and 2,129 objects from the VUDS survey. The redshift distribution of these objects is shown in the top panel of Figure 2.
@@ 2.2.2 Targeted surveys
We also drew on spectra from the C3VO program – the Charting Cluster Construction with VUDS (Le Fèvre et al. 2015) and ORELSE (Lubin et al. 2009) survey – which used the DEIMOS (Faber et al. 2003) and Multi-Object Spectrometer For Infra-Red Exploration (MOSFIRE; McLean et al. 2012) instruments on Keck to follow up protocluster candidates selected with the Voronoi Monte Carlo (VMC) method (Lemaux et al. 2022). Across the fields of the Canada-France-Hawai’i Telescope Legacy Survey D1 (CFHTLS-D1), the Extended Chandra Deep Field South (ECDFS), and COSMOS, C3VO has obtained $\sim\!2000$ high-quality redshifts, yielding detailed maps of structures such as PCl J1000+0200 ($z\sim 2.90$; Cucciati et al. 2014), PCl J0227-0421 ($z\sim 3.31$; Lemaux et al. 2014; Shen et al. 2021), Elentári ($z\sim 3.33$; McConachie et al. 2022; Forrest et al. 2023), Smruti ($z\sim 3.47$; Forrest et al. 2017; Shah et al. 2024), Taralay ($z\sim 4.57$; Lemaux et al. 2018; Staab et al. 2024), and – central to this work – Hyperion.
@@
Additionally, we included data from the Massive Ancient Galaxies at $z>3$ Near-infrared Survey (MAGAZ3NE; Forrest et al. 2024a), a spectroscopic follow-up program that targeted ultra-massive galaxies ($\log_{10}(M_{*}/M_{\odot})\gtrsim 11$ at $z>3$) using Keck/MOSFIRE in multiple fields, including COSMOS. While the galaxies from this survey do not lie within Hyperion, their spectroscopic redshifts provided valuable constraints on photometric redshift outliers. A total of 37 galaxies from MAGAZ3NE met our selection criteria.
Finally, we incorporated spectra from previous studies that targeted individual peaks of Hyperion. Specifically, we included four spectra from Casey et al. (2015), three from Chiang et al. (2015), and one from Diener et al. (2015), all of which met the criteria outlined above. Further details regarding observations and data reduction can be found in the respective publications. The redshift distribution of galaxies from the various targeted spectroscopic surveys is shown in the middle panel of Figure 2.
Figure 2: The distribution of spectroscopic redshifts used in this study from (top) wide-field spectroscopic surveys (section 2.2.1), (middle) targeted spectroscopic surveys (section 2.2.2), and (bottom) the HST-Hyperion survey (section 2.3). For each survey, we show only the galaxies that meet the imposed criteria on R.A., Dec., and IRAC magnitude described in their respective sections. Ground-based spectroscopic redshifts all have reliabilities of $\gtrsim 70\%$, whereas HST-grism redshifts all have reliabilities $\gtrsim 56.7\%$. The number of spectra used from each survey is reported in the legend next to the survey name. In each plot, we additionally mark the approximate boundaries of Hyperion in redshift space.